# Enhancing Scene Graph Generation with Hierarchical Relationships and Commonsense Knowledge

### v2

@@ Abstract
This work introduces an enhanced approach to generating scene graphs by incorporating both a relationship hierarchy and commonsense knowledge. Specifically, we begin by proposing a hierarchical relation head that exploits an informative hierarchical structure. It jointly predicts the relation super-category between object pairs in an image, along with detailed relations under each super-category. Following this, we implement a robust commonsense validation pipeline that harnesses foundation models to critique the results from the scene graph prediction system, removing nonsensical predicates even with a small language-only model. Extensive experiments on Visual Genome and OpenImage V6 datasets demonstrate that the proposed modules can be seamlessly integrated as plug-and-play enhancements to existing scene graph generation algorithms. The results show significant improvements with an extensive set of reasonable predictions beyond dataset annotations. Codes are available at https://github.com/bowen-upenn/scene_graph_commonsense.
@@ 1 Introduction
This work presents simple yet effective approaches in the field of scene graph generation [22, 43, 68, 21, 84, 9]. Scene graph generation, a complex problem that deduces both objects in an image and their pairwise relationships, moves beyond object detection methods [52, 55, 8] which isolate individual object instances. Instead, it represents the entire image as a graph, where each object instance forms a node and the relationships between nodes form directed edges.
@@
Although an advanced scene graph generation model may achieve good performance indicated by its high recall scores [43, 60], it may produce a wide range of unreasonable relationships that are unlikely to occur in the real world, such as ”bunny jumping plate”, even with a high confidence. Recent advances in large language models (LLMs) [6, 1, 63, 61, 5, 20, 4] and vision-language models (VLMs) [49, 41, 40, 53, 66, 44] now enables machines to perform commonsense reasoning [67, 7, 82]. Therefore, we incorporate LLMs or VLMs into the system to critique the output of a scene graph generation model, removing predicates that are not accord with common sense intuitions.
@@ 2 Related Work
ELEGANT [81] leverages LLMs to propose potential relation candidates based on common sense and utilizes BLIP-2 [30], a visual question answering model, for validation. [71] uses CLIP [50] to verify extracted triplets from knowledge bases, and RECODE [32] further aids CLIP with visual cues from LLMs. [45] prompts VLMs to generate scene graphs in JSON formats. VLM4SGG [26] and other works [83, 37, 72] allow language supervisions for weakly supervised generation, and [34, 74, 18, 80] further extend the scope to open vocabularies.
@@ 3 Scene Graph Construction
Figure 2: This diagram provides an overview of HIERCOM. The core components of HIERCOM are the hierarchical relation head and the commonsense validation pipeline, both of which are model-agnostic plug-and-play modules, suitable for integration with a variety of baseline scene graph generation models that have a conventional flat classification head. Specifically, the hierarchical relation head is designed to replace this flat layer, jointly estimating relation super-categories and more granular relations within each category. Additionally, the diagram depicts a baseline scene graph generation model: an RGB image is the input, supplemented by an estimated depth map. Using the DETR object detector, the model generates feature maps, object labels, and bounding boxes. Relationship estimation between each pair of instances occurs in two separate passes - to account for directional relationships - first assuming one instance as the subject and then the other. Subsequently, the commonsense validation pipeline leverages an LLM or VLM - which can have a small size - to filter out commonsense-violating predicates, thus refining the final scene graphs to include more predicates that align with the common sense.
@@ 3.2 Hierarchical relation head
To train the hierarchical relation head, we apply cross-entropy losses to all four terms in Equation 2-5. In addition, we use a supervised contrastive loss [25] to minimize the distances in embedding space within the same relation class (set $P(ij)$) and maximize those from different relationship classes (set $N(ij)$). Except for losses in training DETR [8], the total loss $\mathcal{L}$ can be decomposed as the weighted sum of the following terms:
@@ 3.3 Commonsense validation
Language or vision-language models can critique output predictions of a scene graph generation algorithm and select those that align with common sense. In our settings, we intend to only leverage open-sourced, small-scale language models like LLaMA-3-8B [63] or small vision-language models like LLaVA-1.6-7B [40] to handle this complex graphical task. This choice is motivated by the potential interests in applying these algorithms on local devices, particularly within the robotics community [46, 3, 2, 38], where computations and space are crucial considerations. We also provide a comparison with larger, commercial alternatives like GPT-3.5 [6] in Section 4.
@@ 3.4 Seamless integration with existing frameworks
Figure 2 illustrates how the proposed (1) hierarchical relation head and (2) commonsense validation pipeline can be easily integrated as plug-and-play modules into not only our baseline model but also other existing SOTA scene graph generation algorithms. Specifically, the hierarchical relation head can replace the final linear layer in the classification heads of these models, refining the process of relationship classification. Subsequently, the predicted triplets can undergo the commonsense validation pipeline, eliminating nonsensical ones in the final outputs.
@@ 4 Experiments
NeuralMotifs [76] 58.5 65.2 67.1 32.9 35.8 36.5 21.4 27.2 30.3
NeuralMotifs+HIERCOM (ours) 55.5 69.5 75.6 34.0 41.6 44.8 20.3 28.3 33.9
VTransE [78] 59.0 65.7 67.6 35.4 38.6 39.4 23.0 31.3 35.5
VTransE+HIERCOM (ours) 54.8 68.5 75.0 34.7 41.1 44.3 23.2 31.6 35.9
@@
NeuralMotifs [76] 11.7 14.8 16.1 6.7 8.3 8.8 4.9 6.8 7.9
NeuralMotifs+HIERCOM (ours) 16.3 25.1 30.6 9.8 14.6 17.3 5.7 8.5 10.6
VTransE [78] 11.6 14.7 15.8 6.7 8.2 8.7 3.7 5.0 6.0
VTransE+HIERCOM (ours) 17.9 26.6 32.2 10.3 15.1 17.8 6.7 9.6 12.0
@@ 4.3 Visual results
Figure 4 showcases some predicted scene graphs, which include their top $10$ most confident predicates. The second row displays final scene graphs and the third row shows an ablation without commonsense validation.
The commonsense validation pipeline successfully efficiently reduces many unreasonable predicates that would otherwise have appeared on top of the confidence ranking, for example, it removes the predicate “tree has hand” in the left-most image from its initial prediction. However, in the right-most image, the LLM fails at “shirt on stand”, a possible relation in the real world but unrelated to the scene. Furthermore, it is important to note that there are numerous valid predictions aligned with intuitions that are not annotated; these are marked in black, whose number significantly exceeds the number of pink edges representing true positives according to dataset annotations. This discrepancy highlights the sparse nature of the dataset’s annotations. Besides, all blue edges are incorrect predictions based on both the original dataset annotations and our observations.
Still, our model learns rich relational information even being trained from incomplete annotations. We strongly believe that creating an extensive set of predicates is beneficial for practical scene understanding.
@@ 4.4 More results on the hierarchical relation head
This paragraph discusses an alternative design choice that automatic clusters the relation hierarchy. Instead of adhering to a manually defined hierarchy outlined in Neural Motifs [76], the clustering procedure does not necessarily require human involvement. For example, we can employ k-means, a classic unsupervised clustering technique, on pretrained embedding space to construct the relation hierarchy autonomously. Table 2 shows results using the CLIP-Text [50], GPT-2 [6], and BERT [14] embedding space by transforming relation labels as words into corresponding token embeddings. It turns out that CLIP performs the closest to the manual one in [76] with even higher mR@k scores, so CLIP could be utilized to generalize our idea to new datasets at any scale without the need for manual clustering.
Table 2: This table presents ablation studies on different clustering methods used to structure the relation hierarchy. Results are evaluated on Baseline+HIER, i.e., without commonsense validation for the PredCLS task on Visual Genome. We compare manual clustering defined in [76] that categorizes relations into geometric, possessive, and semantic super-categories, with automatic clustering approaches. The latter ones transform relation labels into CLIP-Text, GPT-2, or BERT’s token embedding space and apply k-means clustering with $k=3$. Results are comparably strong, suggesting that the proposed hierarchical relation head can be effectively generalized to new, larger datasets using automatic clustering methods, thus eliminating the need for manual effort.
@@
Manual [28] 61.1 73.6 78.1 14.4 20.6 23.7
CLIP-Text  [50] 61.6 72.7 76.8 14.4 20.6 23.7
GPT-2 [6] 61.6 69.9 72.0 17.5 25.9 23.7
BERT [14] 61.5 69.7 72.5 16.2 23.0 27.1
Table 3: This table illustrates the integration of the proposed hierarchical relation head with NICE [31] and IETrans [77], two state-of-the-art methods tailored to tackle the long-tailed distribution problem of relation labels. Typically, algorithms addressing long-tailed distributions may compromise R@$k$ scores; however, results from the PredCLS task on Visual Genome demonstrate that incorporating the relation hierarchy not only continues to improve mR@$k$ scores but also enhances their R@$k$ simultaneously.
Methods R@50 R@100 mR@50 mR@100
Motifs+NICE [31] 55.1 57.2 29.9 32.3
Motifs+NICE+HIER (ours) 58.2 65.4 33.1 39.8
Motifs+IETrans [77] 48.6 50.5 35.8 39.1
Motifs+IETrans+HIER (ours) 60.4 66.4 38.0 44.1
Handling the long-tailed distribution
We integrates the proposed modules into scene graph generation algorithms that are specifically tailored to address the long-tailed distribution of relation labels [31, 77]. Numerical results in Table 3 shows that our model-agnostic modules can continue raising their SOTA mR@$k$ scores while simultaneously achieving even higher R@k scores. Furthermore, the histograms in Figure 5 illustrates that we can gain further improvements on the labels at the tail of the distribution.
@@
NeuralMotifs [76] 10.9/14.5 2.2/3.0 0.1/0.2
NeuralMotifs+HIERCOM (ours) 18.4/25.5 4.9/6.4 2.1/3.1
VTransE [78] 11.3/14.7 2.5/3.3 0.8/1.5
VTransE+HIERCOM (ours) 20.1/26.8 5.6/7.3 1.9/3.1
@@
Table 4 details the impact of incorporating depth maps and the supervised contrastive loss in Equation 8 on our baseline model.
@@
Table 7: This table provides a comprehensive ablation study on the choice of foundation models for the commonsense validation pipeline. We show results using LLaMA-3-8B [63], GPT-3.5 [6], and LLaVA-1.6-7B [40] on a variety of baseline models equipped with the hierarchical relation head. The performances across these models are largely comparable, with LLaVA-1.6-7B showing a slight advantage. This finding is particularly promising as it suggests that our commonsense validation pipeline can be effectively deployed using only a small-scale, open-sourced language model on local devices, making it accessible for practical use.
@@ 4.5 More results on the commonsense validation
We implement the commonsense validation pipeline on LLaMA-3-8B [63] and LLaVA-1.6-7B [40] on a single NVIDIA V100 GPU with 32GB memory, and GPT-3.5 [6] through its API calls. Due to high R@$20$ scores in Table 1, we skip the top $m=10$ most confident predictions and validate the next $20$ ones per image. Figure 6 demonstrate examples, including failure cases.
@@
We explore the possibility of baking commonsense knowledge from LLMs into our baseline scene graph generation model. During an initial training phase, we categorize predicates into those that align with commonsense ($\mathcal{S}_{\text{aligned}}$) and those that violate it ($\mathcal{S}_{\text{violated}}$). After collection, we retrain the same model from scratch, but with an additional loss $\mathcal{L}_{\mathrm{cs}}=\mathbf{1}_{\notin\mathcal{S}_{\text{aligned }}}*\lambda_{\text{weak }}+\mathbf{1}_{\in\mathcal{S}_{\text{violated }}}*\lambda_{\text{strong }}$. Here, $\lambda_{\text{weak }}$=$0.1$ and $\lambda_{\text{strong }}$=$10$ serve as penalty weights, and $\mathbf{1}$ is the indicator function. We find only little difference within $1\%$ between querying LLMs at inference and using distilled models. This is encouraging since the latter eliminates the necessity to access LLMs at testing time. We will further explore this idea in the future.

### v3

@@ Abstract
This work introduces an enhanced approach to generating scene graphs by incorporating both a relationship hierarchy and commonsense knowledge. Specifically, we begin by proposing a hierarchical relation head that exploits an informative hierarchical structure. It jointly predicts the relation super-category between object pairs in an image, along with detailed relations under each super-category. Following this, we implement a robust commonsense validation pipeline that harnesses foundation models to critique the results from the scene graph prediction system, removing nonsensical predicates even with a small language-only model. Extensive experiments on Visual Genome and OpenImage V6 datasets demonstrate that the proposed modules can be seamlessly integrated as plug-and-play enhancements to existing scene graph generation algorithms. The results show significant improvements with an extensive set of reasonable predictions beyond dataset annotations. Codes are available at https://github.com/bowen-upenn/scene_graph_commonsense [^1] This work has been accepted at the 2025 IEEE/CVF Winter Conference on Applications of Computer Vision (WACV)..
@@ 1 Introduction
This work presents simple yet effective approaches in the field of scene graph generation [24, 46, 22, 90, 9]. Scene graph generation, a complex problem that deduces both objects in an image and their pairwise relationships, moves beyond object detection or segmentation methods [56, 59, 8, 85, 35] which isolate individual object instances. Instead, it represents the entire image as a graph, where each object instance forms a node and the relationships between nodes form directed edges.
@@
Although an advanced scene graph generation model may achieve good performance indicated by its high recall scores [46, 64], it may produce a wide range of unreasonable relationships that are unlikely to occur in the real world, such as ”bunny jumping plate”, even with a high confidence. Recent advances in large language models (LLMs) [6, 1, 67, 65, 5, 21, 4] and vision-language models (VLMs) [52, 44, 43, 23, 57, 70, 47] now enables machines to perform commonsense reasoning [71, 7, 88]. Therefore, we incorporate LLMs or VLMs into the system to critique the output of a scene graph generation model, removing predicates that are not accord with common sense intuitions.
@@ 2 Related Work
ELEGANT [87] leverages LLMs to propose potential relation candidates based on common sense and validate them using BLIP-2 [32]. [83] collaborates multiple LLMs for enhanced scene understanding and LLM-AR [53] utilizes LLMs as action recognizers. [75] uses CLIP [54] to verify extracted triplets from knowledge bases, and RECODE [34] further aids CLIP with visual cues from LLMs. [48] prompts VLMs to generate scene graphs in JSON formats. VLM4SGG [28] and other works [89, 40, 76] allow language supervisions for weakly supervised generation, and [37, 78, 18, 86] further extend the scope to open vocabularies.
@@ 3 Scene Graph Construction
Figure 2: This diagram provides an overview of HIERCOM. The core components of HIERCOM are the hierarchical relation head and the commonsense validation pipeline, both of which are model-agnostic plug-and-play modules, suitable for integration with a variety of baseline scene graph generation models that have a conventional flat classification head. Specifically, the hierarchical relation head is designed to replace this flat layer, jointly estimating relation super-categories and more granular relations within each category. Additionally, the diagram depicts a baseline scene graph generation model: an RGB image and depth maps are the inputs, and it generates feature maps, object labels, and bounding boxes using DETR. Relations estimation between each pair of instances occurs in two separate passes to account for directional relationships, first assuming one instance as the subject and then the other. Subsequently, the commonsense validation pipeline leverages an LLM or VLM - which can have a small size - to filter out commonsense-violating predicates
@@ 3.2 Hierarchical relation head
To train the hierarchical relation head, we apply cross-entropy losses to all four terms in Equation 2-5. In addition, we use a supervised contrastive loss [27] to minimize the distances in embedding space within the same relation class (set $P(ij)$) and maximize those from different classes (set $N(ij)$). Except for those in DETR [8], the total loss $\mathcal{L}$ can be decomposed as the weighted sum of the following terms:
@@ 3.3 Commonsense validation
Language or vision-language models can critique output predictions of a scene graph generation algorithm and select those that align with common sense. In our settings, we intend to only leverage open-sourced, small-scale language models like LLaMA-3-8B [67] or small vision-language models like LLaVA-1.6-7B [43] to handle this complex graphical task. This choice is motivated by the potential interests in applying these algorithms on local devices, particularly within the robotics community [49, 3, 2, 41], where computations and space are crucial considerations. We also provide a comparison with larger, commercial alternatives like GPT-3.5 [6] in Section 4 with examples in Figure 6.
@@ 3.4 Seamless integration with existing frameworks
Figure 2 illustrates how the proposed HIERCOM - with (1) the “HIER” part: hierarchical relation head and (2) the “COM” part: commonsense validation pipeline - can be easily integrated as plug-and-play modules into not only our baseline model but also other existing SOTA algorithms. Specifically, the hierarchical relation head can replace the final linear layer in the classification heads, refining the process of relationship classification. Subsequently, the predicted triplets can undergo the commonsense validation pipeline, eliminating nonsensical ones in the final outputs.
@@ 4 Experiments
Motifs [80] 58.5 65.2 67.1 32.9 35.8 36.5 21.4 27.2 30.3
Motifs+HIERCOM (ours) 55.5 69.5 75.6 34.0 41.6 44.8 20.3 28.3 33.9
Transformer [63] 59.0 65.7 67.6 35.4 38.6 39.4 23.0 31.3 35.5
Transformer+HIERCOM (ours) 54.8 68.5 75.0 34.7 41.1 44.3 23.2 31.6 35.9
@@
Motifs [80] 11.7 14.8 16.1 6.7 8.3 8.8 4.9 6.8 7.9
Motifs+HIERCOM (ours) 16.3 25.1 30.6 9.8 14.6 17.3 5.7 8.5 10.6
Transformer [63] 11.6 14.7 15.8 6.7 8.2 8.7 3.7 5.0 6.0
Transformer+HIERCOM (ours) 17.9 26.6 32.2 10.3 15.1 17.8 6.7 9.6 12.0
@@ 4.3 Visual results
Figure 4 showcases some predicted scene graphs, which include their top $10$ most confident predicates. The second row displays final scene graphs and the third row shows an ablation without commonsense validation. The commonsense validation successfully efficiently reduces many unreasonable predicates that would otherwise have appeared on top of the confidence ranking, for example, it removes the predicate “tree has hand” in the left-most image from its initial prediction. Furthermore, it is important to note that there are numerous valid predictions aligned with intuitions that are not annotated; these are marked in black, whose number significantly exceeds the number of pink edges representing true positives according to dataset annotations. This discrepancy highlights the sparse nature of the dataset’s annotations. Besides, all blue edges are incorrect predictions based on both the dataset annotations and our observations. Still, our model learns rich relational information even being trained from incomplete annotations. We strongly believe that creating an extensive set of predicates is beneficial for practical scene understanding.
@@ 4.4 More results on the hierarchical relation head
Handling the long-tailed distribution
We integrate the proposed modules into algorithms that are specifically tailored to address the long-tailed distribution of relation labels [33, 81]. We also present the performance of CogTree [77] and HML [13]. Since both methods employ their own hierarchical structures, we did not integrate our hierarchical prediction head to avoid unnecessary complexity. While their methods achieve higher mR@$k$ scores, they exhibit significantly lower R@$k$ scores. Numerical results in Table 2 shows that our model-agnostic modules can continue raising their SOTA mR@$k$ scores while simultaneously achieving even higher R@k scores. Furthermore, the histograms in Figure 5 illustrate that we can gain further improvements on the labels at the tail of the distribution.
Table 2: This table focuses on methods tailored to tackle the long-tailed distribution problem of relation labels. We show results from the PredCLS task. [77, 13, 78] employ their own hierarchical structures, so we did not integrate our hierarchical prediction head to avoid unnecessary redundancy. [77, 13] achieve a different balance with much lower R@$k$ but higher mR@$k$ than HIERCOM on Motifs, Transformer, and VCTree in Table 1. Besides, algorithms addressing long-tailed distributions typically compromise R@$k$ scores significantly; however, results on [33, 81] demonstrate that incorporating the relation hierarchy not only continues to improve mR@$k$ scores but also enhances their R@$k$ simultaneously.
Methods R@50 R@100 mR@50 mR@100
Motifs+CogTree [77] 35.6 36.8 26.4 29.0
Transformer+CogTree [77] 38.4 39.7 28.4 31.0
VCTree+CogTree [77] 44.0 45.4 27.6 29.7
Motifs+HML [13] 47.1 49.1 36.3 38.7
Transformer+HML [13] 45.6 47.8 33.3 35.9
VCTree+HML [13] 47.0 48.8 36.9 39.2
Motifs+CaCao [78] - - 37.1 38.9
Motifs+CaCao+HIERCOM (ours) - - 37.6 39.9
Motifs+NICE [33] 55.1 57.2 29.9 32.3
Motifs+NICE+HIERCOM (ours) 58.2 65.4 33.1 39.8
Motifs+IETrans [81] 48.6 50.5 35.8 39.1
Motifs+IETrans+HIERCOM (ours) 60.4 66.4 38.0 44.1
@@
We discuss an alternative design choice that automatically clusters the relation hierarchy. Instead of adhering to a manually defined hierarchy outlined in Neural Motifs [80], the clustering procedure does not necessarily require human involvement. For example, we can employ k-means, a classic unsupervised clustering technique, on pretrained embedding space to construct the relation hierarchy autonomously. Table 3 shows results using the CLIP-Text [54], GPT-2 [6], and BERT [14] embedding space by transforming relation labels as words into corresponding token embeddings. The commonsense validation is also seamlessly integrated here. It turns out that CLIP performs the closest to the manual one in [80] with even higher mR@k scores, so CLIP could be utilized to generalize our idea to new datasets at any scale without the need for manual clustering.
Table 3: This table presents ablation studies on different clustering methods used to structure the relation hierarchy. Results are evaluated on Baseline+HIERCOME. We compare manual clustering defined in [80] that categorizes relations into geometric, possessive, and semantic super-categories, with automatic clustering approaches. The latter ones transform relation labels into CLIP-Text, GPT-2, or BERT’s token embedding space and apply k-means clustering with $k=3$. Results are comparably strong, suggesting that the proposed hierarchical relation head can be effectively generalized to new, larger datasets using automatic clustering methods, thus eliminating the need for manual effort.
@@
Manual [30] 64.2 75.6 79.2 17.7 23.9 26.7
CLIP-Text  [54] 61.9 73.6 77.4 17.1 25.6 30.5
GPT-2 [6] 62.5 70.5 72.7 17.9 25.0 29.7
BERT [14] 61.8 69.9 72.8 16.3 23.3 27.4
@@
Motifs [80] 10.9/14.5 2.2/3.0 0.1/0.2
Motifs+HIERCOM (ours) 18.4/25.5 4.9/6.4 2.1/3.1
Transformer [82] 11.3/14.7 2.5/3.3 0.8/1.5
Transformer+HIERCOM (ours) 20.1/26.8 5.6/7.3 1.9/3.1
@@
The first row of each section in Table 7 shows the ablation with the relation hierarchy but without the commonsense validation. Table 4 details the impact of incorporating depth maps and the supervised contrastive loss in Equation 8 on our baseline model.
@@
Table 7: This table provides a comprehensive ablation study on (1) models with hierarchical relation head but without commonsense validation, i.e., “+HIER” in the first row of each section, and (2) the choice of foundation models for commonsense validatio, including LLaMA-3-8B [67], GPT-3.5 [6], and LLaVA-1.6-7B [43]. The performances across these models are largely comparable, with LLaVA-1.6-7B showing a slight advantage. This finding is particularly promising as it suggests that our commonsense validation pipeline can be effectively deployed using only a small-scale, open-sourced language model on local devices, making it accessible for practical use.
@@ 4.5 More results on the commonsense validation
We implement the commonsense validation pipeline on LLaMA-3-8B [67] and LLaVA-1.6-7B [43] on a single NVIDIA V100 GPU with 32GB memory (around 3.37s and 4.75s per image on Motifs+HIERCOM), and GPT-3.5 [6] through its API calls. We skip the top $m=10$ most confident predictions and validate the next $20$ ones per image. To minimize LLM’s hallucinations [19], we exclude validating any subject-relation-object combinations in the training dataset, resulting in $0\%$ unique false-negative triplets in training and $0.27\%$ in testing.
@@
To address slowdowns in inference caused by incorporating LLMs, we explore the option of embedding commonsense knowledge from LLMs directly into our baseline model. During an initial training phase, we categorize predicates into those that align with commonsense ($\mathcal{S}_{\text{aligned}}$) and those that violate it ($\mathcal{S}_{\text{violated}}$). After this classification, we can retrain the same model from scratch but with an additional loss $\mathcal{L}_{\mathrm{cs}}=\mathbf{1}_{\notin\mathcal{S}_{\text{aligned }}}*\lambda_{\text{weak }}+\mathbf{1}_{\in\mathcal{S}_{\text{violated }}}*\lambda_{\text{strong }}$. Here, $\lambda_{\text{weak }}$=$0.1$ and $\lambda_{\text{strong }}$=$10$ serve as penalty weights, and $\mathbf{1}$ is an indicator. We find only little difference within $1\%$ between querying LLMs at inference and using distilled models. Alternative to distillation, we can also filter out predicates based on $\mathcal{S}_{\text{violated}}$ during future runs of inference. Both options are encouraging since they eliminates the necessity to access LLMs at test time.